# The dual of Philo’s shortest line segment problem

### v2

@@ Abstract.
We study the dual of Philo’s shortest line segment problem which asks to find the optimal line segments passing through two given points, with a common endpoint, and with the other endpoints on a given line. The provided solution uses multivariable calculus and geometry methods. Interesting connections with the angle bisector of the triangle are explored. A similar property of a symedian of a triangle is conjectured.
@@ 1. Introduction
Maximum-minimum problems in geometry played crucial role in the history of development of mathematical methods [11], Chapter VII, [8], [21] Chapters VIII-XIX. Many attempts to solve variety of geometry problems for optimization lead to the develoment of differential and integral calculus and later to the calculus of variations. Maxima-minima problems also play important role in the pedagogy of mathematics [3], [25]. In mathematical olympiads for schools or universities, and journals with problem solving columns such as Crux Mathematicorum and American Mathematical Monthly one can often find at least one geometric inequality included in its list of proposed problems. Recent breakthough in automation of proofs for classical euclidean geometry theorems is yet to be done for the geometric inequalities [27]. Mathematicians noticed very early that some of the geometric optimization problems do not have solutions and even when they do have a solution it is not always possible to construct it with euclidean instruments (a compass and a straightedge). The problem about minimal line segment with fixed point and inscribed into a given angle is one of them [9], p. 405-406, [23], p. 36-37, [24] p. 33, 179-181, [18], [5], [19]. In this problem an angle $ABC$ and a point $D$ in it is given and it is asked to find the line $EF$ through this point which has shortest line segment $EF$ cut by this angle. Sir I. Newton mentioned this problem as an example for his method of fluxions (see [20], p. 45, Problem 5.III). The solution of this problem is known in the literature as Philo’s line [13], [14] p. 39 and 234-238, [29] p. 182-183, [28], [16] p. 115–116. If $BG$ is the perpendicular to $EF$ from point $B$, then $EF$ is shortest when $|ED|=|FG|$ (see Figure 1).
@@
Philo or Philon of Byzantium found its connection with the classical problem of the dublication of cube [7], p. 107, [31], [22] p. 24, [26]. The digits of the length of Philo’s line through the incenter of triangle with sidelengths 3,4,5 appear in the OEIS as A195284 [17]. Generalizations for non-euclidean geometry and higher dimensions appeared in [12], [30]. Related problem of the longest or shortest chords of conics and curves in general have similar solutions [4], [6], [2].
In projective geometry the principle of duality is used to explain the similarity between theorems where the roles of lines and points are interchanged. If we use this idea in the above extremum problem, then we obtain the following dual problem ([2], p. 4). Given line $l$ and points $A$ and $B$ find point $C$ with ray $CA$ and $CB$ intersecting line $l$ at $D$ and $E$, respectively, such that the sum of lengths of $CD$ and $CE$ is minimal (see Figure 1 and Table 1). This problem is surprisingly much more challenging than Philo’s shortest line problem. In the following text this problem is solved using combination of methods from calculus and geometry. At the end of the paper we mentioned other interesting formulae about the configuration and several open problems.
@@
As was mentioned before the optimal solution of geometric extrema problems can sometimes be non-constructible with the help of a compass, a ruler, and the given data. Without the proof of the existence of the optimal solution the proof is incomplete. For example, in Hadamard’s classical geometry problem book some of such extrema problems are provided with a comment like "assume that a maximum exists" (see e.g. Exercise 366 in p. 313, Exercise 418b in p. 322 of [15]). Most of the time it is possible to refer to calculus results for more rigorous proof but it makes the solution less appealing for geometrically inclined mind. Another possible way out from this situation is to give or construct the optimal solution first in the diagram and then ask to prove that the other possibilities are not optimal. In the problem about Philo’s line we can give $\triangle BEF$ and points $D$ and $G$ on its side $EF$ such that $|ED|=|FG|$, $BG\perp EF$, and then prove that any other line segment $E_{1}F_{1}$ through point $D$ has a greater length (see Figure 1). We can formulate the mentioned dual problem above in the same way by giving the optimal solution first. It will be clear from the proof why it is optimal and how it was found in the first place.
@@
Let $CDE$ be a triangle with acute angles at vertices $D$ and $E$. Let $CK$ be the angle bisector of triangle $CDE$. Drop perpendiculars $KL$ and $KM$ to sides $CD$ and $CE$, respectively. Take points $A$ and $B$ on sides $CD$ and $CE$, respectively, such that $|AC|=|LD|$, $|BC|=|ME|$. Then for any point $C_{1}$ different from $C$, such that rays $C_{1}A$ and $C_{1}B$ intersect line $DE$ at points $D_{1}$ and $E_{1}$, respectively, inequality $|C_{1}D_{1}|+|C_{1}E_{1}|>|CD|+|CE|$ holds true (see Figure 2).
@@
Let us denote the coordinates of points as $A(a,b),\ B(c,d),\ D_{1}(x_{1},0)$ and $E_{1}(x_{2},0)$. Without loss of generality we can assume that $a<c$ and $b\geq d>0$. If we denote coordinates of $C_{1}$as $(x,y)$, then $x_{1}=\frac{ay-bx}{y-b}$ and $x_{2}=\frac{cy-dx}{y-d}$. Here $(x,y)\in\mathbb{D}$, where $\mathbb{D}=\{(x,y)|x\in(-\infty,+\infty),\ y>b\}$. The last inequality is needed to guarantee the condition in the satatement of the problem which says that rays $C_{1}A$ and $C_{1}B$ intersect line $DE$. We obtain that
@@
Note that $f(x,y)$ is continuous in $\mathbb{D}$ and $f(x,y)\rightarrow+\infty$ as $x\rightarrow\pm\infty$, $y\rightarrow b$ when $x\neq a$. Similarly, $f(x,y)\rightarrow+\infty$ as $y\rightarrow+\infty$. Also $\lim_{(x,y)\rightarrow(a,b)}f(x,y)$ does not exist and depends on the direction of approach to point $(a,b)$. Indeed, if $b>d$, then
@@
where $k_{2}>0$. This limit is minimal when $k_{1}=0$ which corresponds to the direction of line $x=a$. If $b=d$, then the last limit is also infinite. Function $f(x,y)$ is also bounded from below because it is a positive function. Then either there is a point $(x^{*},y^{*})\in\mathbb{D}$ where function $f(x,y)$ is minimal or this function approaches its infimum when $(x,y)\rightarrow(a,b)$ $(y>b)$. We will study the last case later. In the first case $f^{\prime}_{x}(x^{*},y^{*})=0$, $f^{\prime}_{y}(x^{*},y^{*})=0$. Obviously, $a\leq x^{*}\leq c$, because if $x<a$, then $f(x,y)>f(a,y)$, and if $x>c$, then $f(x,y)>f(c,y)$. Note also that if $y$ is fixed, then function $f(x,y)$ is convex with respect to $x$. Indeed,
@@
$f_{x}(a,y)=\frac{y\left(a-c\right)}{\left(y-d\right)\sqrt{{\left(y-d\right)^{2}+\left(a-c\right)^{2}}}}<0,$
$f_{x}(c,y)=\frac{y\left(c-a\right)}{\left(y-b\right)\sqrt{{\left(y-b\right)^{2}+\left(a-c\right)^{2}}}}>0.$
Therefore $a<x^{*}<c$. In order to locate point $x^{*}$, note that equality $f^{\prime}_{x}(x,y)=0$ implies
@@
The last equality can be interpreted geometrically in the following way. Take points $L_{1}$ and $M_{1}$ on sides $C_{1}D_{1}$ and $C_{1}E_{1}$, respectively, such that $|AC_{1}|=|L_{1}D_{1}|$, $|BC_{1}|=|M_{1}E_{1}|$. Take points $K_{1}$ and $K_{2}$ on line $D_{1}E_{1}$ such that $K_{1}L_{1}\perp C_{1}D_{1}$ and $K_{2}M_{1}\perp C_{1}E_{1}$. Equality $f^{\prime}_{x}(x,y)=0$ is equivalent to say that $|K_{1}L_{1}|=|K_{2}M_{1}|$ (see Figure 3). Indeed, if we draw the line through $C_{1}$ and parallel to $D_{1}E_{1}$ and take points $F$ and $G$ such that $AF\perp AC_{1}$ and $BG\perp BC_{1}$, then $|AF|=|K_{1}L_{1}|$ and $|BG|=|K_{2}M_{1}|$. On the other hand,
@@
The other condition ($f^{\prime}_{y}(x,y)=0$) can also be interpreted geometrically but its little harder than the previous condition. First, drop perpendicular $C_{1}H$ to side $D_{1}E_{1}$. Next, we calculate
$f^{\prime}_{y}(x,y)=\left(\sqrt{\left(\frac{x-a}{y-b}\right)^{2}+1}+\sqrt{\left(\frac{x-c}{y-d}\right)^{2}+1}\right)+$
@@
If $a<x<c$, then equality $f^{\prime}_{y}(x,y)=0$ can be interpreted as
@@
from which it follows that (2) holds true. Since $\frac{|D_{1}H|^{2}}{|L_{1}D_{1}|}=|D_{1}H|\cdot\frac{|C_{1}D_{1}|}{|D_{1}K_{1}|}$ and $\frac{|E_{1}H|^{2}}{|M_{1}E_{1}|}=|E_{1}H|\cdot\frac{|C_{1}E_{1}|}{|E_{1}K_{2}|}$, equality (2) can also be written as
@@
If $AU$ is the angle bisector of triangle $D_{1}C_{1}E_{1}$, then by the angle bisector theorem $\frac{|D_{1}C_{1}|}{|C_{1}E_{1}|}=\frac{|D_{1}U|}{|UE_{1}|}$. By noting this we can write (4) as
@@
If we denote $|D_{1}H|=v$, $|K_{2}H|=w$, $|K_{2}U|=z$, $|K_{1}U|=t$, $|K_{1}E_{1}|=u$ (see Figure 4), equation (5) can be written simply as
@@
It is possible only when $z=t=0$, which means that points $K_{1},K_{2},$ and $U$ coincide. The other cases when points $K_{1}$ or $K_{2}$ are outside of the side $D_{1}E_{1}$, or when point $K_{1}$ is closer to point $D_{1}$ than point $K_{2}$ can be considered similarly. For example, if the order of points is as in Figure 5, where the notations for the lengths of the line segments has slightly changed, then
@@
Finally, if the order of points is as in Figure 6, where again the notations for the lengths of the line segments has slightly changed, then
@@
Note that because of equality (1), we only need to consider the cases when point $U$ is between points $K_{1}$ and $K_{2}$. Otherwise, if $U$ is not between points $K_{1}$ and $K_{2}$, then one can drop perpendiculars from point $U$ to sides $C_{1}D_{1}$ and $C_{1}E_{1}$ and one of them will be larger than the other one, which is impossible. Similarly, we do not need to consider the cases when points $K_{1}$ and $K_{2}$ are on different sides of point $H$, because in these cases one of the sides of equality (3) is positive while the other side is negative and therefore (3) does not hold.
If $\angle D_{1}=\frac{\pi}{2}$ ($x=a$), then point $K_{1}$ goes to infinity and equation (5) simplifies to $\frac{|D_{1}U|}{|UE_{1}|}\cdot\frac{|E_{1}K_{2}|}{|K_{2}H|}=1,$ which simply means that points $K_{2}$ and $U$ coincide. Similarly, if $\angle E_{1}=\frac{\pi}{2}$ ($x=c$), then point $K_{2}$ goes to infinity and equation (5) simplifies to $\frac{|D_{1}U|}{|UE_{1}|}\cdot\frac{|HK_{1}|}{|K_{1}D_{1}|}=1,$ which implies that points $K_{1}$ and $U$ coincide.
We proved that if $C_{1}(x^{*},y^{*})$ is a minimum point of $f(x,y)$, then endpoint $U$ of angle bisector $C_{1}U$ coincides with points $K_{1},K_{2}$. Note that triangle $CDE$ and its angle bisector $CK$ satisfy all these conditions together with (1) and (2). In order to complete the proof, we need to show that there are no minimum points other than $C$.
We will now study the locus of intersection point of line $L_{1}K_{1}$ and $M_{1}K_{2}$ when $|AD_{1}|=|BE_{1}|$ and $D_{1}(x_{1},0)$ changes on line $DE$, and then intersection of this locus with line $DE$ (see Figure 7). If $b=d$, then by symmetry this locus is the perpedicular bisector of the line segment $AB$, which always intersect line $DE$ at a unique point. So, we can assume that $b>d$. Let us denote $|AD_{1}|=|BE_{1}|=t$. Then $x_{1}=a-\sqrt{t^{2}-b^{2}}$ and $x_{2}=c+\sqrt{t^{2}-d^{2}}$. Then the equations of lines $AD_{1}$ and $BE_{1}$ are $\frac{y-b}{x-a}=\frac{-b}{x_{1}-a}$ and $\frac{y-d}{x-c}=\frac{-d}{x_{2}-c}$, respectively. The coordinates of intersection point $C_{1}(x_{0},y_{0})$ of these lines are
@@
The coordinates of point $L_{1}(x_{3},y_{3})$ are $x_{3}=x_{0}+x_{1}-a$, $y_{3}=y_{0}-b$. Similarly, the coordinates of point $M_{1}(x_{4},y_{4})$ are $x_{4}=x_{0}+x_{2}-c$, $y_{3}=y_{0}-d$. Then the equations of lines $L_{1}K_{1}$ and $M_{1}K_{2}$ are $\frac{y-y_{3}}{x-x_{3}}=\frac{x_{1}-a}{b}$ and $\frac{y-y_{4}}{x-x_{4}}=\frac{x_{2}-c}{d}$, respectively. By solving the last two equations for $y$ we calculate ordinate $y^{\prime}$ of intersection point of lines $L_{1}K_{1}$ and $M_{1}K_{2}$ as
@@
where the denominator is always positive and the numerator is a decreasing function for $t\geq b$. The abscissa $x^{\prime}$ of this intersection point can also be determined but we will not need it:
@@
The formula for $x^{\prime}$ and $y^{\prime}$ determine a parametric curve which is locus of intersection point of lines $L_{1}K_{1}$ and $M_{1}K_{2}$. This locus can intersect line $DE$ at maximum 1 point for $t\geq b$. Indeed, note that $\phi(b)=d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}$ and $\phi(+\infty)=-\infty$. Therefore, if $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}\geq 0$, then equation $\phi(t)=0$ has a unique solution in interval $t\geq b$. Otherwise, if $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}<0$, then equation $\phi(t)=0$ has no solutions in interval $t\geq b$. In the statement of the problem we are given acute angled triangle $CDE$ satisfying all the conditions for extrema. Therefore, $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$ and the only solution of equation $\phi(t)=0$ in interval $t>b$ is $t_{0}=|AD|$. Note that $t_{0}^{2}=d^{2}+\lambda^{2}$, where $\lambda$ is one of the solutions of fourth degree equation
@@
$\lim_{y\rightarrow b^{+}}f_{y}(a,y)=\left(1+\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}\right)\left(1-\frac{d}{b-d}\left(\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}-1\right)\right),$
@@
$\frac{b-d\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}}{b-d}=\frac{b(b-d)-d\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}}{\left(b-d\right)^{2}}$
$=\frac{(b^{2}-d^{2})(b-d)^{2}-d^{2}(a-c)^{2}}{(b-d)^{2}\left(b(b-d)+d\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}\right)},$
@@
which is negative because as was mentioned before $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$. Therefore $f(a,y)$ is decreasing and then increasing for $y>b$ and consequently, $f(a,y)$ can not have infimum at $(a,b)$ and the only minimum of $f(x,y)$ is point $C$. ∎

### v7

@@ Abstract.
We study the dual of Philo’s shortest line segment problem and find the optimal line segments passing through two given points, with a common endpoint, and with the other endpoints on a given line. This problem is dual, in a point-and-line-exchanging sense, to a famous problem of antiquity used to solve the problem of duplicating the cube. The provided solution uses multivariable calculus and elementary geometry methods. Interesting connections with the angle bisector of the triangle are explored. A generalization of the problem using $l_{p}$ ($p\geq 1$) norm is proposed. The particular case $p=\infty$ is also studied. It is shown that in the cases $p=0$ and $p=2$ the median and the symedian, respectively, of a triangle do not always give a solution for the corresponding optimization problems. The general case $p\neq 1$ and related problems are proposed as open problems.
@@ 1. Introduction
Maximum-minimum problems in geometry played crucial role in the history of development of mathematical methods [14], Chapter VII, [11], [30] Chapters VIII-XIX. Many attempts to solve variety of geometry problems for optimization lead to the develoment of the differential and integral calculus and later to the calculus of variations. Geometric extrema problems play a crucial role in mathematics education and research [6], [36], Chs. 4 and 13; [44]; [24]; [15], Sect 1.8. In the mathematical olympiads for schools or universities, and the journals with problem solving columns such as Mathematics Magazine, The American Mathematical Monthly, and Crux Mathematicorum one can often find at least one geometric inequality included in its list of proposed problems. The recent breakthough in automation of proofs for classical euclidean geometry theorems is yet to be done for the geometric inequalities [38]. The connections of the geometric extrema problems with $l_{p}$ norm (see for example [23], p. 20) in functional analysis were discussed in [33], [43]. Historically, mathematicians recognized that certain geometric optimization problems lack solutions and even when they do have a solution it is not always possible to construct it with the euclidean instruments (i.e. a pair of compasses and a straightedge). The problem about the minimal line segment with a fixed point and inscribed into a given angle is one of them [12], p. 405-406, [32], p. 36-37, [35] p. 33, 179-181, [26], [8], [27]. In this problem an angle $ABC$ and a point $D$ in it is given and it is asked to find the line $EF$ through this point which has shortest line segment $EF$ cut by this angle (see Appendix). Sir I. Newton mentioned this problem and its generalizations as an example for his method of fluxions (see [29], p. 45, Problem 5.III). The solution of this problem is known in the literature as Philo’s line [17], [18] p. 39 and 234-238, [40] p. 182-183, [39], [21] p. 115–116. If $BG$ is the perpendicular to $EF$ from point $B$, then $EF$ is shortest when $|ED|=|FG|$ (see Figure 1).
@@
Philo or Philon of Byzantium found the solution of the classical problem of the duplication of a cube using this line [10], p. 107, [42], [31] p. 24, [37]. The digits of the length of Philo’s line through the incenter of triangle with sidelengths 3,4,5 appear in the OEIS as A195284 [22]. Generalizations for the non-euclidean geometry and the higher dimensions appeared in [16], [41]. Related problem of the longest or shortest chords of conics and curves in general have similar solutions [7], [9], [2].
There are examples of the use of the principle of duality in metric geometry [25] but the duality is usually used in projective geometry to explain the similarity between the theorems where the roles of lines and points are interchanged. If we use the analogous idea in the above extremum problem, then we obtain the following dual problem in a point-and-line-exchanging sense ([2], p. 4). Given the line $l$ and the points $A$ and $B$ find the point $C$ with the rays $CA$ and $CB$ intersecting the line $l$ at the points $D$ and $E$, respectively, such that the sum of the lengths of the line segments $CD$ and $CE$ is minimal (see Figure 1 and Table 1). Note that the line segments $AC$ and $BC$ themselves do not intersect the line $l$ or equivalently $C$ is on the same open half plane as $A$ and $B$. Furthermore, the points $A$ and $B$ are located between $l$ and $C$. The problem presents a significantly greater challenge compared to Philo’s shortest line problem. The problem is solved using a combination of multivariable calculus and elementary geometric methods. The paper concludes with noteworthy formulas related to the configuration and some open problems for further studies.
@@
As was mentioned before the optimal solution of the geometric extrema problems can sometimes be non-constructible with the help of a pair of compasses, a ruler, and the given data. Assuming existence of optimal configuration without a proper proof can sometimes lead to contradictory results (see [34], Sect. 18a in the original and Sect. 21 in later editions). See also for example [23], p. 56-57 where it is shown how assuming existence of a triangle with maximal sum of angles in non-euclidean geometry leads to a "proof" of the fifth postulate of Euclid. Therefore, without a justification of the existence of the optimal solution the proof is incomplete (cf. [28], Chs. 4 and 12). For example, in Hadamard’s classical geometry problem book some of such extrema problems are provided with a comment like "assume that a maximum exists" (see e.g. Exercise 366 in p. 313, Exercise 418b in p. 322 of [20]). Most of the time it is possible to refer to calculus results for more rigorous proof with compact sets and Weierstrass theorem but it makes the solution less appealing from the geometric point of view. Another possible way out from this situation is to give or construct the optimal solution first in the diagram and then ask to prove that the other possibilities are not optimal. In [19], Sect. 1.3 these two approaches to geometric extrema problems are referred to as indirect and direct proofs, respectively (cf. [35], pp. 7-10). In the problem about Philo’s line the direct proof would be giving $\triangle BEF$ and points $D$ and $G$ on the side $EF$ such that $|ED|=|FG|$, $BG\perp EF$, and then proving that any other line segment $E_{1}F_{1}$ through point $D$ has a greater length (see Appendix). We can formulate the mentioned dual problem above in the same way by giving the optimal solution first. The proof establishes the optimality of the solution and its derivation.
@@
Let $CDE$ be a triangle with acute angles at the vertices $D$ and $E$. Let $CK$ be the angle bisector of the triangle $CDE$. Drop the perpendiculars $KL$ and $KM$ to the sides $CD$ and $CE$, respectively. Take the points $A$ and $B$ on the sides $CD$ and $CE$, respectively, such that $|AC|=|LD|$, $|BC|=|ME|$. Then for any point $C_{1}$ different from $C$, such that the rays $C_{1}A$ and $C_{1}B$ intersect the line $DE$ at the points $D_{1}$ and $E_{1}$, respectively, inequality $|C_{1}D_{1}|+|C_{1}E_{1}|>|CD|+|CE|$ holds true (see Figure 2).
@@
Let us denote the coordinates of the points as $A(a,b),\ B(c,d),\ D_{1}(x_{1},0)$ and $E_{1}(x_{2},0)$. Without loss of generality we can assume that $a<c$ and $b\geq d>0$. If we denote coordinates of $C_{1}$as $(x,y)$, then $x_{1}=\frac{ay-bx}{y-b}$ and $x_{2}=\frac{cy-dx}{y-d}$. Here $(x,y)\in\mathbb{D}$, where $\mathbb{D}=\{(x,y)|x\in(-\infty,+\infty),\ y>b\}$. The last inequality is needed to guarantee the condition in the statement of the problem which says that the rays $C_{1}A$ and $C_{1}B$ intersect the line $DE$. We obtain that
@@
Note that the function $f(x,y)$ is continuous in $\mathbb{D}$ and $f(x,y)\rightarrow+\infty$ as $x\rightarrow\pm\infty$, $y\rightarrow b$ when $x\neq a$. Similarly, $f(x,y)\rightarrow+\infty$ as $y\rightarrow+\infty$. Also $\lim_{(x,y)\rightarrow(a,b)}f(x,y)$ does not exist and depends on the direction of approach to the point $(a,b)$. Indeed, if $b>d$, then
@@
where $k_{2}>0$. This limit is minimal when $k_{1}=0$ which corresponds to the direction of the line $x=a$. If $b=d$, then the last limit is also infinite. Function $f(x,y)$ is also bounded from below because it is a positive function. Then either there is a point $(x^{*},y^{*})\in\mathbb{D}$ where function $f(x,y)$ is minimal or this function approaches its infimum when $(x,y)\rightarrow(a,b)$ $(y>b)$. We will study the last case later. In the first case $f_{x}(x^{*},y^{*})=0$, $f_{y}(x^{*},y^{*})=0$. It is not hard to compute that $f_{x}$ is negaitve if $x<a$ and positive if $x>c$. Therefore, if $x<a$, then $f(x,y)>f(a,y)$, and if $x>c$, then $f(x,y)>f(c,y)$, and consequently $a\leq x^{*}\leq c$. Note also that if $y$ is fixed, then the function $f(x,y)$ is convex with respect to $x$. Indeed,
@@
$f_{x}(a,y)=\frac{y\left(a-c\right)}{\left(y-d\right)\sqrt{{\left(a-c\right)^{2}+\left(y-d\right)^{2}}}}<0,$
$f_{x}(c,y)=\frac{y\left(c-a\right)}{\left(y-b\right)\sqrt{{\left(a-c\right)^{2}+\left(y-b\right)^{2}}}}>0.$
Therefore $a<x^{*}<c$. In order to locate the point $x^{*}$, note that equality $f_{x}(x,y)=0$ implies
@@
The last equality can be interpreted geometrically in the following way. Take the points $L_{1}$ and $M_{1}$ on the sides $C_{1}D_{1}$ and $C_{1}E_{1}$, respectively, such that $|AC_{1}|=|L_{1}D_{1}|$, $|BC_{1}|=|M_{1}E_{1}|$. Take the points $K_{1}$ and $K_{2}$ on the line $D_{1}E_{1}$ such that $K_{1}L_{1}\perp C_{1}D_{1}$ and $K_{2}M_{1}\perp C_{1}E_{1}$. The equality $f^{\prime}_{x}(x,y)=0$ is equivalent to say that $|K_{1}L_{1}|=|K_{2}M_{1}|$ (see Figure 3). Indeed, if we draw the line through $C_{1}$ and parallel to $D_{1}E_{1}$ and take the points $F$ and $G$ such that $AF\perp AC_{1}$ and $BG\perp BC_{1}$, then $|AF|=|K_{1}L_{1}|$ and $|BG|=|K_{2}M_{1}|$. On the other hand,
@@
The other condition ($f_{y}(x,y)=0$) can also be interpreted geometrically but its little harder than the previous condition. First, drop the perpendicular $C_{1}H$ to the side $D_{1}E_{1}$. Next, we calculate
$f_{y}(x,y)=\left(\sqrt{\left(\frac{x-a}{y-b}\right)^{2}+1}+\sqrt{\left(\frac{x-c}{y-d}\right)^{2}+1}\right)+$
@@
If $a<x<c$, then the equality $f_{y}(x,y)=0$ can be interpreted as
@@
from which it follows that (2) holds true. Since $\frac{|D_{1}H|^{2}}{|L_{1}D_{1}|}=|D_{1}H|\cdot\frac{|C_{1}D_{1}|}{|D_{1}K_{1}|}$ and $\frac{|E_{1}H|^{2}}{|M_{1}E_{1}|}=|E_{1}H|\cdot\frac{|C_{1}E_{1}|}{|E_{1}K_{2}|}$, the equality (2) can also be written as
@@
If $AU$ is the angle bisector of the triangle $D_{1}C_{1}E_{1}$, then by the angle bisector theorem $\frac{|D_{1}C_{1}|}{|C_{1}E_{1}|}=\frac{|D_{1}U|}{|UE_{1}|}$. By noting this we can write (4) as
@@
In the following the proof will break into three cases (Figures 4, 5, and 6) depending on the relative positions of $K_{1}$ and $K_{2}$. If we denote $|D_{1}H|=v$, $|K_{2}H|=w$, $|K_{2}U|=z$, $|K_{1}U|=t$, $|K_{1}E_{1}|=u$ (see Figure 4), equation (5) can be written simply as
@@
It is possible only when $z=t=0$, which means that the points $K_{1},\ K_{2},$ and $U$ coincide. The other cases when the points $K_{1}$ or $K_{2}$ are outside of the side $D_{1}E_{1}$, or when the point $K_{1}$ is closer to the point $D_{1}$ than the point $K_{2}$ can be considered similarly. For example, if the order of the points is as in Figure 5, where the notations for the lengths of the line segments has slightly changed, then
@@
Finally, if the order of the points is as in Figure 6, where again the notations for the lengths of the line segments has slightly changed, then
@@
Note that because of the equality (1), we only need to consider the cases when the point $U$ is between the points $K_{1}$ and $K_{2}$. Otherwise, if $U$ is not between the points $K_{1}$ and $K_{2}$, then one can drop perpendiculars from the point $U$ to the sides $C_{1}D_{1}$ and $C_{1}E_{1}$ and one of them will be larger than the other one, which is impossible. Similarly, we do not need to consider the cases when the points $K_{1}$ and $K_{2}$ are on the different sides of point $H$, because in these cases one of the sides of the equality (3) is positive while the other side is negative and therefore (3) does not hold.
If $\angle D_{1}=\frac{\pi}{2}$ ($x=a$), then the point $K_{1}$ goes to infinity and the equation (5) simplifies to $\frac{|D_{1}U|}{|UE_{1}|}\cdot\frac{|E_{1}K_{2}|}{|K_{2}H|}=1,$ which simply means that the points $K_{2}$ and $U$ coincide. Similarly, if $\angle E_{1}=\frac{\pi}{2}$ ($x=c$), then the point $K_{2}$ goes to infinity and the equation (5) simplifies to $\frac{|D_{1}U|}{|UE_{1}|}\cdot\frac{|HK_{1}|}{|K_{1}D_{1}|}=1,$ which implies that the points $K_{1}$ and $U$ coincide.
[…]
First, note that if $CK$ is the angle bisector, and $KL$ and $KM$ are the perpendiculars to the sides (see Fig. 2), then $CL$ and $CM$ are congruent. Hence, if $AC=LD$ and $BC=ME$, then $AD=BE$. We will use this fact to study the locus of the intersection point of lines $L_{1}K_{1}$ and $M_{1}K_{2}$ when $|AD_{1}|=|BE_{1}|$ and $D_{1}(x_{1},0)$ changes on line $DE$, and then intersection of this locus with line $DE$ (see Figure 7). If $b=d$, then by the symmetry this locus is the perpedicular bisector of the line segment $AB$, which always intersect line $DE$ at a unique point. So, we can assume that $b>d$. Let us denote $|AD_{1}|=|BE_{1}|=t$. Then $x_{1}=a-\sqrt{t^{2}-b^{2}}$ and $x_{2}=c+\sqrt{t^{2}-d^{2}}$. Then the equations of the lines $AD_{1}$ and $BE_{1}$ are $\frac{y-b}{x-a}=\frac{-b}{x_{1}-a}$ and $\frac{y-d}{x-c}=\frac{-d}{x_{2}-c}$, respectively. The coordinates of the intersection point $C_{1}(x_{0},y_{0})$ of these lines are
@@
The coordinates of the point $L_{1}(x_{3},y_{3})$ are $x_{3}=x_{0}+x_{1}-a$, $y_{3}=y_{0}-b$. Similarly, the coordinates of the point $M_{1}(x_{4},y_{4})$ are $x_{4}=x_{0}+x_{2}-c$, $y_{3}=y_{0}-d$. Then the equations of the lines $L_{1}K_{1}$ and $M_{1}K_{2}$ are $\frac{y-y_{3}}{x-x_{3}}=\frac{x_{1}-a}{b}$ and $\frac{y-y_{4}}{x-x_{4}}=\frac{x_{2}-c}{d}$, respectively. By solving the last two equations for $y$ we calculate the ordinate $y^{\prime}$ of intersection point of lines $L_{1}K_{1}$ and $M_{1}K_{2}$ as
@@
where the denominator is always positive and the numerator is a decreasing function for $t\geq b$. It is because in the numerator of $\phi(t)$ the first term, which is negative, is increasing in absolute value and the second term, which is positive, is decreasing.
The abscissa $x^{\prime}$ of this intersection point can also be determined but we will not need it:
@@
The formula for $x^{\prime}$ and $y^{\prime}$ determine a parametric curve which is the locus of intersection point of lines $L_{1}K_{1}$ and $M_{1}K_{2}$. This locus can intersect line $DE$ at maximum 1 point for $t\geq b$. Indeed, note that $\phi(b)=d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}$ and $\phi(+\infty)=-\infty$. Therefore, if $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}\geq 0$, then equation $\phi(t)=0$ has a unique solution in interval $t\geq b$. Otherwise, if $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}<0$, then equation $\phi(t)=0$ has no solutions in interval $t\geq b$. In the statement of the problem we are given acute angled triangle $CDE$ satisfying all the conditions for extrema. Therefore, $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$ and the only solution of equation $\phi(t)=0$ in interval $t>b$ is $t_{0}=|AD|$. Note that $t_{0}^{2}=d^{2}+\lambda^{2}$, where $\lambda$ is one of the solutions of fourth degree equation
@@
$\lim_{y\rightarrow b^{+}}f_{y}(a,y)=\left(1+\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}\right)\left(1-\frac{b}{b-d}\frac{\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}-1}{\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}}\right),$
@@
$\frac{b(b-d)-d\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}}{\left(b-d\right)\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}}$
$=\frac{(b^{2}-d^{2})(b-d)^{2}-d^{2}(a-c)^{2}}{(b-d)\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}\left(b(b-d)+d\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}\right)},$
@@
which is negative because as was mentioned before $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$. Therefore $f(a,y)$ is decreasing and then increasing for $y>b$ and consequently, $f(a,y)$ can not have infimum at $(a,b)$ and the only minimum of $f(x,y)$ is the point $C$. ∎